NEXT
Make New Entry,
Make Followup Entry
User name riordan
Log entry time 18:13:52 on March 09, 2009
Entry number 265480
keyword=Summary for March 8 Swing - March 9 Day
Summary for Mar 8 swing - Mar 9 Day
Finished 4G starting 4E.
Vacuum levels become too great and the machine goes down at 2:30
Beam is breifly restored around 9:00. MCC spends the day warming up
1L04, which will not be restored for this running.
BigBite and LHRS are switched to positive polarity
Beam in Use
Shift |
ABU |
BANU |
BNA |
2009/03/08 Swing | 4.71 |
2.47 | 0.82 |
2009/03/09 Owl | 2.14 |
0.03 | 5.83 |
2009/03/09 Day | 0.48 |
3.52 | 4.01 |
TOTAL | 7.33 |
6.02 | 10.66 |
Main events/issues:
- Mar 8
- Magnet cryo problems reappeared around 8:00pm. Called MCC and
the cryo engineer found that the bottom fill valve for Left Q2 was
closed. Problem was corrected and has not reappeared.
- BHWP IN
- Mar 9
- 02:30 beam goes away due to vacuum problems
- 08:30 beam is restored briefly
- 10:20 Hall is on controlled access for magnet polarity change
- 12:00 MCC decides to warm up 1L04. No beam for some time.
Plans for the next few days:
- Positive polarity running on BigBite and LHRS when beam is restored
3/10 owl
- At 8:00AM Tuesday, decisions will be made if the machine is to be
reconfigured based on early morning running.
- Continue 4-pass transverse production
- Compton electron detector material installation is delayed.
A copy of this log entry has been emailed to: brads, jpchen, jiang, meziani, choi, riordan